If an assignment has 40 question And is worth 12 points and six are wrong what is the grade?
Answer by nerdybill(7384)

If an assignment has 40 question And is worth 12 points and six are wrong what is the grade?
total questions answered correctly:
40-6 = 34
Grade:
points awarded = 34*12 = 408
total possible points = 40*12 = 480
.
Grade = 408/480 * 100 = 0.85 * 100 = 85%